# To Offload or Not To Offload: Model-driven Comparison of Edge-native and On-device Processing

### v2

@@ III-C1 Comparing Edge Offload and On-Device Execution
Lemma III.1.
For accelerator-driven workloads, edge offloading incurs a higher average end-to-end latency than on-device processing when {dmath} s_dev - s_edge ¡ λμdevnetdevnet- λ) + λμedgenetedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc )
@@
$w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$ (3)
Let $D_{req}$ and $D_{res}$ denote the task and result payload size, respectively. Further, let $B$ denote the bandwidth of the device-to-edge network path. The above inequality reduces to
$s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$ (4)
As shown in [29], DNN inference on programmable accelerators often exhibits deterministic service times (denoted as $D$), as the number of operations per request remains constant. Since typical accelerators have a small degree of parallelism $k$ [30], the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process. While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ [31, 32]. This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$. Its expected wait time can be derived from the Pollaczek-Khinchin (P-K) formula [33, 19] and is given by
$E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$ (5)
@@
$E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$ (6)
Let $k_{dev}$ and $k_{edge}$ denote the degree of parallelism seen at the accelerator on the device and the edge respectively. Applying these two results to (4), we get {dmath} s_dev - s_edge ¡ 1μdevnet- λdevnet - 1μdevnet + 1μedgenet- λedgenet - 1μedgenet + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λedge - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
Offloading is only required when a request arrives, which gives us $\lambda^{net}_{dev}=\lambda$. For a queuing system to be stable, throughput (i.e., task completion rate) must equal the task arrival rate $\lambda$ under steady-state conditions. When tasks are offloaded to an edge server, results are transmitted to the device only upon task completion, which yields $\lambda_{edge}^{net}=\lambda$. The first two terms involving $\mu_{dev}^{net}$ on the RHS can be written as {dmath} 1μdevnet- λdevnet - 1μdevnet = λμdevnet(μdevnet- λ) whereas the second two terms involving $\lambda_{edge}^{net}$ can be written as {dmath} 1μedgenet- λedgenet - 1μedgenet = λμedgenet(μedgenet- λ).
Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$. Substituting the above into (III-C1), we obtain {dmath} s_dev - s_edge ¡ λμdevnet(μdevnet- λ) + λμedgenet(μedgenet- λ) + Dreq+ DresB + 12 ( 1kedgeμedgeproc- λ - 1kedgeμedgeproc ) - 12 ( 1kdevμdevproc- λ - 1kdevμdevproc ) which completes the proof. ∎
@@
Explanation. Service time, which is the time spent processing a request on an accelerator, is determined by the computational demand of the workload divided by the processing capacity of the hardware. For any given hardware configuration, requests with lighter demand have to be processed much faster at the edge to overcome the network penalty for edge offloading to be faster than on-device processing. For requests with lower computational demands, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally due to the disparity in processing capacities between the device and the edge as the workload computational demand decreases. Further, since the service rate is the inverse of service times in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right hand side of the inequality in (III.1) will also become smaller since $\mu$ increases for lighter requests. The terms representing the network overheads remain unchanged since they depend on request size rather than the processing demand of the request. As a result, the inequality in (III.1) is more likely to hold true for requests with lighter processing demand. ∎
Remark III.2.
@@
Explanation. The term $\frac{D_{req}+D_{res}}{B}$ in (III.1) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right hand size of the inequality. In addition, longer transmission times reduce the service rates of the NIC at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the RHS. As a result, the inequality is more likely to hold for slower networks or tasks with higher network overheads, making on-device processing faster than edge offloading. ∎
@@
Practical takeaways. Lemma III.1 establishes a bound on the average service time difference between device and edge for edge offloading to be more effective. The bound and the resulting inequality depends on several factors: (i) relative processing speeds of the device and edge, denoted by $s_{dev}$, $s_{edge}$ (and also the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) network speed, (iii) overall workload denoted by request rate $\lambda$, (iv) network transmission overheads. In general, for any specific workload, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading the workload; otherwise, on-device processing becomes preferable to offloading.
@@
Moreover, the request rate $\lambda$ impacts offloading decisions in two competing ways. First, as $\lambda$ increases, both the edge and the device experience longer queuing delays, with the edge queue growing more slowly due to its faster processing speed. Consequently, the benefits of offloading improve under high load. However, under slow network conditions or for applications with large payloads, a high $\lambda$ increases the network queuing delay (first two terms of the right side of the inequality) and may lead to network congestion, making on-device processing preferable. Thus, the optimal strategy depends on the workload’s payload size relative to both network bandwidth and the edge-device performance gap.
@@
Remark III.1 shows how the computational demand of the workload influences the execution strategy. For workloads with higher computational demand, edge offloading is more likely to outperform on-device processing, as the edge’s processing speed advantage becomes larger. Conversely, for workloads with lower computational demand, on-device processing is more efficient as the performance advantage of edge processing is likely outweighed by the associated network costs. Remark III.2 demonstrates how network delays affect the efficiency of edge offloading. With smaller network bandwidth or larger request/result payloads, both the network queuing delay and transmission time increase. As a result, edge offloading becomes less efficient, while on-device processing is unaffected since it does not rely on network resources.
@@
Extending to collaborative processing. Collaborative (i.e., split) processing involves partially processing requests locally on the client’s device before offloading them to a remote edge server to minimize the amount of data sent over the network [34, 35, 36, 12, 26, 37, 38]. To compute the end-to-end response time of such a strategy, we combine our analytic models for on-device processing and edge offloading into a tandem queuing network that accounts for both the partial local computation and the subsequent transmission and processing delays at the edge server. Specifically, a request first enters the components modeled in Figure 1b and is partially processed with a service time $s^{\prime}_{dev}$ representing the partial local processing time. Then, the intermediate output of size $D_{inter}$ passes through the components modeled in Figure 1a, with a network transmission time of $D_{inter}/B$ and a service time $s^{\prime}_{edge}$ representing the remaining computation time on the edge. This combined model can then be used to estimate performance and inform decision-making when selecting between local, edge, or collaborative processing strategies.
III-D Impact of Multi-Tenancy
Figure 2: $m$ devices offloading requests to an edge server.
@@ III-D Impact of Multi-Tenancy
While the analysis in Lemma III.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work. Hence, we extend our model-driven comparison to multi-tenancy scenarios.
@@
Modeling edge application multiplexing. In the case where each edge server is dedicated to a single client device, the workload (i.e., the request arrival rate) at the edge and the device $\lambda^{proc}_{dev}$ and $\lambda^{proc}_{edge}$ are identical. However, when an edge server is multiplexed across $m$ devices, its workload will be the sum of the individual device workloads under the offloading case, as shown in Figure 2. This higher workload will increase the queuing delays at the edge and may cause the end-to-end offloading latency to exceed the on-device processing latency.
@@
Since the request arrivals at each client device are governed by independent Poisson process, and the composition of independent Poisson processes results in a new Poisson process [39], it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ that is the sum of the device workload. That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the $i-th$ device.
@@
We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary. For this case, we model each system as an $M/G/1$ system, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the device and the edge. This extension captures the potentially large variability in service times across applications in multi-tenant environments. If $s_{i}$ denotes the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times. That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ and the the the effective service rate of the edge $\mu_{edge}=1/s_{edge}$. Then the aggregate utilization of the edge server $\rho_{edge}$ is $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$. This leads us to the following lemma for the multi-tenant case.
Lemma III.2.
@@
For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when {dmath} s_dev - s_edge ¡ λdevμdevnetdevnet- λdev) + λedgeμedgenetedgenet- λedge) + Dreq+ DresB + ρedge+ λedgekedgeμedgeprocVar[sedge]2(kedgeμedgeproc- λedge) - 12 ( 1kdevμdevproc- λdev - 1kdevμdevproc )
@@
The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by the Pollaczek-Khinchin (P-K) formula [33, 19] {dmath} E[w] = ρ+ λμVar[s]2(μ- λ) Substituting (III-D) for $w_{edge}^{proc}$, (5) for $w_{dev}^{proc}$, and (6) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into (4) completes the proof. ∎
@@
Practical takeaways. The expected wait time in $M/G/1/FCFS$ systems is heavily affected by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays. It is therefore important for application designers to consider service time variability when assigning applications to edge servers. Grouping applications with similar processing demands can help reduce overall variability, thereby minimizing queuing delays and improving edge offloading performance. Further, the above analysis assumes that isolation features such as GPU virtualization, which are available on server-grade GPUs, are not available on edge accelerators. If such features were available, each device could be assigned to an isolated virtual GPU on the edge. In this special case, the offloading situation is reduced to a situation similar to that considered in Sec III-C1 where each virtual GPU presents $\frac{1}{m}$-th of the capacity to each device. However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
III-E Discussion

### v3

@@ 3.3.1. Comparing Edge Offload and On-Device Execution
Lemma 3.0.
For accelerator-driven workloads, edge offloading incurs a higher average end-to-end latency than on-device processing when
(3) $s_{dev}-s_{edge}<\frac{\lambda}{\mu_{dev}^{net}(\mu_{dev}^{net}-\lambda)}+\frac{\lambda}{\mu_{edge}^{net}(\mu_{edge}^{net}-\lambda)}+\frac{D_{req}+D_{res}}{B}+\frac{1}{2}\left(\frac{1}{k_{edge}\mu_{edge}^{proc}-\lambda}-\frac{1}{k_{edge}\mu_{edge}^{proc}}\right)-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
@@
(4) $w_{dev}^{proc}+s_{dev}<w_{dev}^{net}+n_{req}+w_{edge}^{proc}+s_{edge}+w_{edge}^{net}+n_{res}$
Let $D_{req}$ and $D_{res}$ denote the request and result payload size, respectively. Further, let $B$ denote the bandwidth of the device-to-edge network path. The above inequality reduces to
(5) $s_{dev}-s_{edge}<w_{dev}^{net}+w_{edge}^{net}+\frac{D_{req}}{B}+\frac{D_{res}}{B}+w_{edge}^{proc}-w_{dev}^{proc}$
As shown in (Liang2023), DNN inference on accelerators exhibits deterministic service times (denoted as $D$), as the number of operations per request remains constant. Since typical accelerators have a small degree of parallelism $k$ (Amert2017), the queuing system can be modeled as an $M/D/k/FCFS$ system, where request arrivals follow a Poisson process. While no closed-form solution exists for the expected wait time in such a system, a common approximation is to aggregate the service rate across the degree of parallelism $k$ (Shah2018; Shah2020). This reduces the model to an $M/D/1/FCFS$ system with aggregated service rate $k\mu$. Its expected wait time can be derived from the Pollaczek-Khinchin (P-K) formula (Gross2008; Harchol2013) and is given by
(6) $E[w_{M/D/1}]=\frac{1}{2}\left(\frac{1}{k\mu-\lambda}-\frac{1}{k\mu}\right)$
@@
(7) $E[w_{M/M/1}]=\frac{1}{\mu-\lambda}-\frac{1}{\mu}$
Let $k_{dev}$ and $k_{edge}$ denote the degree of parallelism seen at the accelerator on the device and the edge, respectively. Applying these two results to (5), we get
(8) $s_{dev}-s_{edge}<\frac{1}{\mu_{dev}^{net}-\lambda_{dev}^{net}}-\frac{1}{\mu_{dev}^{net}}+\frac{1}{\mu_{edge}^{net}-\lambda_{edge}^{net}}-\frac{1}{\mu_{edge}^{net}}+\frac{D_{req}+D_{res}}{B}+\frac{1}{2}\left(\frac{1}{k_{edge}\mu_{edge}^{proc}-\lambda_{edge}}-\frac{1}{k_{edge}\mu_{edge}^{proc}}\right)-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda_{dev}}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
Offloading is required only when a request arrives, which gives us $\lambda^{net}_{dev}=\lambda$. For a queuing system to be stable, throughput (i.e., task completion rate) must equal the task arrival rate $\lambda$ under steady-state conditions. When tasks are offloaded to an edge server, results are transmitted to the device only upon task completion, which yields $\lambda_{edge}^{net}=\lambda$. Moreover, regardless of the execution strategy, the application sees the same arrival rate, i.e., $\lambda_{dev}^{proc}=\lambda_{edge}^{proc}=\lambda$. Substituting the above into (8), we obtain
(9) $s_{dev}-s_{edge}<\frac{\lambda}{\mu_{dev}^{net}(\mu_{dev}^{net}-\lambda)}+\frac{\lambda}{\mu_{edge}^{net}(\mu_{edge}^{net}-\lambda)}+\frac{D_{req}+D_{res}}{B}+\frac{1}{2}\left(\frac{1}{k_{edge}\mu_{edge}^{proc}-\lambda}-\frac{1}{k_{edge}\mu_{edge}^{proc}}\right)-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
which completes the proof. ∎
@@
Explanation. Service time, the time spent processing a request on an accelerator, is determined by the workload’s computational demand divided by the hardware’s processing capacity. For any given hardware configuration, requests with lower computational demands have to be processed much faster at the edge to overcome the network penalty for edge offloading to outperform on-device processing. For requests with lower demand, the difference in average service times $s_{dev}-s_{edge}$ shrinks proportionally as the disparity in processing capacity between the device and the edge becomes less impactful. Further, since the service rate is the inverse of service time in queuing theory (i.e., $s=1/\mu$), the terms with $\mu_{dev}^{proc}$ and $\mu_{edge}^{proc}$ on the right-hand side of the inequality in (3) will also become smaller since $\mu$ increases for lighter requests. Meanwhile, the terms representing the network overheads remain unchanged, as they depend on the payload size rather than the request’s processing demand. As a result, the inequality in (3) is more likely to hold for requests with lighter processing demand. ∎
Remark 3.3.
@@
Explanation. The term $\frac{D_{req}+D_{res}}{B}$ in (3) increases when either the network bandwidth $B$ decreases or the request and result sizes $D_{req}$ and $D_{res}$ become larger, thereby increasing the right-hand side of the inequality. In addition, longer transmission times reduce the NIC service rates at the device and edge ($\mu_{dev}^{net}$ and $\mu_{edge}^{net}$), further increasing the right-hand side. As a result, the inequality is more likely to hold under slower networks or for tasks with larger payloads, making on-device processing faster than edge offloading. ∎
@@
Practical takeaways. Lemma 3.1 establishes a bound on the average service time difference between the device and the edge for edge offloading to be more effective. The bound and the resulting inequality depend on several factors: (i) relative processing speeds of the device and the edge $s_{dev}$ and $s_{edge}$ (and the corresponding service rates $\mu_{dev}$ and $\mu_{edge}$), (ii) overall workload denoted by request rate $\lambda$, (iii) network transmission overheads. In general, the edge processing capacity must be larger than the device processing capacity by a factor that overcomes the network penalty of offloading; otherwise, on-device processing becomes preferable to offloading.
@@
Moreover, the request rate $\lambda$ impacts offloading decisions in two competing ways. First, as $\lambda$ increases, both the edge and the device experience longer queuing delays, though the edge queue grows more slowly due to its faster processing speed, which favors offloading. Second, under slow networks or for tasks with large payloads, a high $\lambda$ increases network queuing delay, which penalizes offloading. Consequently, the optimal strategy depends on the workload’s payload size relative to network bandwidth and the performance gap between the device and the edge.
@@
Remark 3.2 indicates that edge offloading is likely to outperform on-device processing for workloads with high computational demand as the edge’s processing speed advantage dominates, while on-device processing is more efficient for lighter workloads since network overhead can outweigh the edge’s advantage. Remark 3.3 shows that edge offloading becomes less efficient with limited bandwidth or large payloads, which increase network queuing and transmission delays, whereas on-device processing remains unaffected.
@@
Extending to collaborative processing. Collaborative (i.e., split) processing involves partially processing requests locally on the device before offloading them to a remote edge server to minimize the amount of data sent over the network (Almeida2022; Banitalebi2021; Hu2019; Huang2020; Kang2017; Li2018; Yang2022; Chen2025). To compute the end-to-end response time of such a strategy, we combine our analytic models for on-device processing and edge offloading into a tandem queuing network that accounts for both the partial local computation and the subsequent transmission and processing delays at the edge server. Specifically, a request first enters the components modeled in Figure 1b and is partially processed with a service time $s^{\prime}_{dev}$ representing the partial local processing time. Then, the intermediate output of size $D_{inter}$ passes through the components modeled in Figure 1a, with a network transmission time of $D_{inter}/B$ and a service time $s^{\prime}_{edge}$ representing the remaining computation time on the edge. This combined model can then be used to estimate performance and inform decision-making when selecting between local, edge, or collaborative processing.
3.4. Impact of Multi-Tenancy
@@
While the analysis in Lemma 3.1 implicitly assumes an edge server that is dedicated to a client device, in practice, edge servers will be multiplexed across multiple client devices and service multiple applications that offload their work. Hence, we extend our model-driven comparison to multi-tenancy scenarios.
@@
Modeling edge application multiplexing. In the case where each edge server is dedicated to a single client device, the workload (i.e., the request arrival rate) at the edge and at the device $\lambda^{proc}_{dev}$ and $\lambda^{proc}_{edge}$ are identical. However, when an edge server is multiplexed across $m$ devices, its workload will be the sum of the individual device workloads under the offloading case, as shown below:
@@
This higher workload will increase the queuing delays at the edge and may cause the end-to-end offloading latency to exceed the on-device processing latency. Since the request arrivals at each client device are governed by independent Poisson processes, and the composition of independent Poisson processes results in a new Poisson process (pishro2014), it follows that the aggregate edge workload is also Poisson with an arrival rate $\lambda_{edge}$ equal to the sum of the workloads of all devices. That is, $\lambda_{edge}=\sum_{i=0}^{m}\lambda_{i}.$, where $\lambda_{i}$ denotes the workload at the ithi-th device.
@@
We assume that the multiplexed applications have different service time demands, and the aggregate service time distribution seen at the edge will be arbitrary. In this case, we model the edge system as an $M/G/1$ system, where $M$ denotes Poisson arrivals and $G$ denotes the general (arbitrary) service time distribution seen at the edge. This extension captures the potentially large variability in service times across applications in multi-tenant environments. Let $s_{i}$ denote the average service time of requests of device $i$ when processed at the edge, then $s_{edge}$ is the weighted average of these service times. That is, $s_{edge}=\sum_{i}\frac{\lambda_{i}}{\lambda_{edge}}s_{i}$ with the effective service rate of the edge given by $\mu_{edge}^{proc}=1/s_{edge}$. The aggregate utilization of the edge server is then $\rho_{edge}=\lambda_{edge}/\mu_{edge}^{proc}$. This leads us to the following lemma for the multi-tenant case.
Lemma 3.0.
@@
For multi-tenant edge servers, the average end-to-end latency of edge offloading is higher than on-device processing when
(10) $s_{dev}-s_{edge}<\frac{\lambda_{dev}}{\mu_{dev}^{net}(\mu_{dev}^{net}-\lambda_{dev})}+\frac{\lambda_{edge}}{\mu_{edge}^{net}(\mu_{edge}^{net}-\lambda_{edge})}+\frac{D_{req}+D_{res}}{B}+\frac{\rho_{edge}+\lambda_{edge}k_{edge}\mu_{edge}^{proc}\text{Var}[s_{edge}]}{2(k_{edge}\mu_{edge}^{proc}-\lambda_{edge})}-\frac{1}{2}\left(\frac{1}{k_{dev}\mu_{dev}^{proc}-\lambda_{dev}}-\frac{1}{k_{dev}\mu_{dev}^{proc}}\right)$
@@
The expected wait time of $M/G/1/FCFS$ systems is well-known and is given by P-K formula (Gross2008; Harchol2013)
(11) $E[w_{M/G/1}]=\frac{\rho+\lambda\mu\text{Var}[s]}{2(\mu-\lambda)}$
Substituting (11) for $w_{edge}^{proc}$, (6) for $w_{dev}^{proc}$, and (7) for $w_{dev}^{net}$ and $w_{edge}^{net}$ into (5) completes the proof. ∎
@@
Practical takeaways. The expected wait time in $M/G/1/FCFS$ systems is heavily influenced by the variance in service times, meaning that co-located applications with highly variable service demands can significantly increase queuing delays. It is therefore important for application designers to consider service time variability when assigning applications to edge servers. Grouping applications with similar processing demands can help reduce overall variability, thereby minimizing queuing delays and improving edge offloading performance. Further, the above analysis assumes that isolation features such as GPU virtualization, which are available on server-grade GPUs, are not available on edge accelerators. If such features were available, each device could be assigned to an isolated virtual GPU on the edge. In this special case, the offloading situation reduces to the one similar to that considered in Sec 3.3.1 where each virtual GPU provides $\frac{1}{m}$-th of the capacity to each device. However, many edge accelerators lack such features, resulting in a shared accelerator that services a combined workload.
3.5. Discussion